# Continual Lie algebras determined by chain complexes

### v4

@@ 4. Conclusion
In conclusion, we would like to mention a few directions of development and further applications of the material presented in this paper. We propose a way to associate a continual Lie algebra to a chain complex. Thus the properties, in particular, Jacobi identity, kernels, and relations of resulting continual Lie algebras depend on the kind of chain complex spaces as well as on the set of their parameters. One can think of introducing various types of products suitable for the construction of systems of differential relations more complicated than (2.2). In our particular case (Section 2), in order to make connection with continual Lie algebras, we have chosen the commutator (2.17) (with respect to the original product defined on bicomplex spaces) as the simplest natural product. One could think of other possibilities which would be coherent with the orthogonality condition (2.5).
@@
The example of the Čhech-de Rham cohomology that we study in Section 3 comes from the differential geometry of foliations. In classics, the orthogonality condition applied to elements and their differentials of one particular bicomplex space, boils down to the integrability condition, and leads to the Frobenius theorem. Then it delivers the Godbillon-Vey cohomological class whose geometric meaning is not yet completely studied [9]. As for further applications in differential geometry, starting from the orthogonality condition, it would be interesting to find other cohomological invariants related to the Čech-de Rham bicomplex for foliations, and to understand their geometric meaning. The constructions considered in this paper can be also used for the cohomology theory of smooth manifolds, in particular, in various approaches to the construction of cohomological classes (cf., in particular, [13]). Wide applications are awaiting for new examples of continual Lie algebras in the field of integrable models [15, 14]. The cases of non-commutative fields used to define continual Lie algebras would also be useful for in non-commutative geometry.
@@ Acknowledgment
5. Appendix: Čech-de Rham complex for foliations
In this Appendix we recall [5] the notion of the basis of transversal sections, and Čech-de Rham complex for a foliation of a smooth manifold. Let $M$ be a smooth manifold of dimension $n$, equipped with a foliation $\mathcal{F}$ of co-dimension $l$ [5]. A transversal section of $\mathcal{F}$ is an embedded $l$-dimensional submanifold $U\subset M$ which is everywhere transverse to $\mathcal{F}$ leaves. If $\alpha$ is a path between two points $x$ and $y$ on the same leaf, and if $U$ and $V$ are transversal sections through $x$ and $y$, then $\alpha$ defines a transport along the leaves from a neighborhood of $x$ in $U$ to a neighborhood of $y$ in $V$. Therefore, we define a germ of a diffeomorphism $hol(\alpha):(U,x)\longrightarrow(V,y)$, called the holonomy of the path $\alpha$. If such a transport is defined in all of $U$ and embeds $U$ into $V$, this embedding $h:U\hookrightarrow V$ is sometimes also denoted by $hol(\alpha):U\hookrightarrow V$. Embeddings of this form will be called holonomy embeddings.
The product is defined by

### v5

@@ 4. Conclusion
In conclusion, we would like to mention a few directions of development and further applications of the material presented in this paper. We propose a way to associate a continual Lie algebra to a chain complex. Thus the properties, in particular, the Jacobi identity, kernels, and relations of resulting continual Lie algebras depend on the kind of chain complex spaces as well as on the set of their parameters. One can think of introducing various types of products suitable for the construction of systems of differential relations more complicated than (2.2). In our particular case (Section 2), in order to make connection with continual Lie algebras, we have chosen the commutator (2.17) (with respect to the original product defined on bicomplex spaces) as the simplest natural product. One could think of other possibilities which would be coherent with the orthogonality condition (2.5).
@@
The example of the Čhech-de Rham cohomology that we study in Section 3 comes from the differential geometry of foliations. In classics, the orthogonality condition applied to elements and their differentials of one particular bicomplex space, boils down to the integrability condition, and leads to the Frobenius theorem. Then it delivers the Godbillon-Vey cohomology class whose geometric meaning is not yet completely studied [9]. As for further applications in differential geometry, starting from the orthogonality condition, it would be interesting to find other cohomology invariants related to the Čech-de Rham bicomplex for foliations, and to understand their geometric meaning. The constructions considered in this paper can be also used for the cohomology theory of smooth manifolds, in particular, in various approaches to the construction of cohomology classes (cf., in particular, [13]). Wide applications are awaiting for new examples of continual Lie algebras in the field of integrable models [15, 14]. The cases of non-commutative fields used to define continual Lie algebras would also be useful for in non-commutative geometry.